OpenStudy (anonymous):

which value of x is a solution of 5/x = x=13/6 A) -2 B) -3 C) -10 D) -15

jimthompson5910 (jim_thompson5910):

too many equal signs are present...

OpenStudy (anonymous):

o sorry i meant x+13/6

jimthompson5910 (jim_thompson5910):

5/x = (x+13)/6 5*6 = x*(x+13) 30 = x^2 + 13x 0 = x^2 + 13x - 30 x^2 + 13x - 30 = 0 (x+15)(x-2) = 0 x + 15 = 0 or x -2 = 0 x = -15 or x = 2 So the solutions are x = -15 or x = 2